I don't understand how to create an algebraic formula to answer this question:
A solution of 75% vinegar is to be mixed with a solution of 22% vinegar to form 636 liters of a 70% solution. How many liters of the 75% solution must be used?

Answer by stanbon(75887)

A solution of 75% vinegar is to be mixed with a solution of 22% vinegar to form 636 liters of a 70% solution. How many liters of the 75% solution must be used?
------
Equation:
vinegar + vinegar = vinegar
-----
0.75x + 0.22(636-x) = 0.70*636
-----
Multiply thru by 100 to get:
75x + 22*636 - 22x = 70*636
-----
53x = 48*636
x = 576 liters (amt. of 75% solution needed)
----
636-x = 60 liters (amt. of 22% solution needed)
====================================================
Cheers,
Stan H.
==================